OpenStudy (jmartinez638):

As a box slides down a ramp, friction does 23.0J of work. At the bottom of the ramp, the box has 3.8J of kinetic energy and is traveling at 2.8 m/sec. How high is the ramp?

OpenStudy (jmartinez638):

@SolomonZelman

OpenStudy (jmartinez638):

Weird problem...

OpenStudy (jmartinez638):

I just do not have any formulas for this, and need some help setting it up

OpenStudy (johnweldon1993):

Yeah working through it now, just dont want to give it away lol

OpenStudy (jmartinez638):

O, thanks!

OpenStudy (johnweldon1993):

Hmm it's weird, I'm missing 1 variable...gimme a sec :) Maybe typing it out here will help

OpenStudy (johnweldon1993):

So first, lets look at the fact they gave us the kinetic energy at the bottom of the ramp 3.8J when traveling at 2.8m/s Remember our equation for kinetic energy? \[\large K.E = \frac{1}{2}mv^2\] Since we are given K.E and 'v' we can solve here for the mass of the box...what would that be?

OpenStudy (johnweldon1993):

You there?

OpenStudy (johnweldon1993):

Regardless, this is just a conservation of energy problem, the potential energy at the top is equal to the kinetic energy at the bottom plus any work done by friction \[\large P.E = K.E + W_f\] After calculating your mass, we write out this equation *with the fact we are given both the kinetic energy and the frictional work \[\large mgh = 3.8J + 23J\] Solving for 'h' we have \[\large h = \frac{26.8J}{mg}\] Plug in your found mass, multiply it by gravity and carry out this equation to find your height!

OpenStudy (jmartinez638):

Ok, sounds good. I'll show my work in a sec!

OpenStudy (jmartinez638):

Thanks, @johnweldon1993

OpenStudy (johnweldon1993):

No problem!

OpenStudy (jmartinez638):

@johnweldon1993 I'm sorry, right as you offered this great explanation, I became occupied with something else, and am just now getting back to this. Solving Kinetic Energy for mass gives me a number in Newton meters. What am I doing wrong?

OpenStudy (jmartinez638):

I understand what is going on, but I am messing up on that specific calculation

OpenStudy (jmartinez638):

Newton*meter is not a unit of mass, but of work done.

OpenStudy (johnweldon1993):

Sorry, was away for a bit So the kinetic energy is the problem here? \[\large K.E = \frac{1}{2}\text{mass}v^2\] \[\large 3.8J = \frac{1}{2}\text{mass}(2.8(\frac{m}{s})^2)\] \[\large 3.8J = \text{mass}(3.92\frac{m^2}{s^2})\] \[\large \text{mass} = \frac{0.97J}{\frac{m^2}{s^2}}\] We know \(\large J = kg\frac{m^2}{s^2}\) so we have \[\large \frac{0.97kg\cancel{\frac{m^2}{s^2}}}{\cancel{\frac{m^2}{s^2}}}\] Leaving \(\large 0.97kg\)

OpenStudy (jmartinez638):

Ah, I see it now. Thanks so much

OpenStudy (jmartinez638):

And the height of the ramp would be ~2.82 meters high.

OpenStudy (jmartinez638):

Sound about right??

OpenStudy (johnweldon1993):

Sorry again >.< yeah that's about what I remember getting!